Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A miniature roulette wheel is divided into 10 equal sectors

Expert replies
by BTGmoderatorLU » Wed Jan 24, 2018 5:24 am
A miniature roulette is divided into 10 equal sectors, each bearing a distinct integer from 1 to 10, inclusive. Each time the wheel is spun, a ball randomly detemines the winning sector by setting in that sector. If the wheel is spun three times, approximately what is the probability that the product of the three winning sectors' integers will be even?

A. 88%
B. 75%
C. 67%
D. 63%
E. 50%

The OA is A.

I'm really confused with this PS question. Experts, any suggestion about how can I solve this question? Thanks in advance.
Join the discussion
Source: — Problem Solving |

by [email protected] » Wed Jan 24, 2018 7:28 pm
Hi LUANDATO,

We're told that a roulette wheel is divided into 10 equal sectors, with each sector bearing a distinct integer from 1 to 10, inclusive. Each time the wheel is spun, a ball randomly detemines the winning sector by setting in that sector. The wheel is spun three times and we're asked for the approximate probability that the PRODUCT of the three winning sectors' integers will be EVEN. This question is based on some Number Properties and Probability math.

To start, the product of any integer and an EVEN number is EVEN. Thus, if any of the 3 numbers is EVEN, then the product will be EVEN. Rather than calculate all of the different outcomes that involve at least one even number, we can instead calculate what we DON'T WANT to have happen (the product is ODD); we can then subtract that probability from the number 1 to determine the probability of what we DO want (the product is EVEN).

With the integers 1-10 inclusive, we have 10 total values (5 odds and 5 evens). The probability of the three numbers all being ODD is....

(Odd)(Odd)(Odd) = (5/10)(5/10)(5/10) = (1/2)(1/2)(1/2) = 1/8

Thus, the probability of AT LEAST one of the three numbers being EVEN is 1 - 1/8 = 7/8

1/8 = .125 = 12.5%
7/8 = .875 = 87.5%

The approximate value of our calculation is 88%

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Jeff@TargetTestPrep » Fri Jan 26, 2018 12:26 pm
LUANDATO wrote:A miniature roulette is divided into 10 equal sectors, each bearing a distinct integer from 1 to 10, inclusive. Each time the wheel is spun, a ball randomly detemines the winning sector by setting in that sector. If the wheel is spun three times, approximately what is the probability that the product of the three winning sectors' integers will be even?

A. 88%
B. 75%
C. 67%
D. 63%
E. 50%
We can use the formula:

P(even product) = 1 - P(odd product)

To get an odd product we must have all odd numbers, to the probability of 3 odds is:

1/2 x 1/2 x 1/2 = 1/8.

So P(even product) = 1 - 1/8 = 7/8 ≈ 88%

Answer: A

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion